OpenStudy (anonymous):

rewrite the slope-intercept form equation into standard form. y=-2x+4

OpenStudy (anonymous):

@sk8king14

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

standard form is Ax+By = C

jimthompson5910 (jim_thompson5910):

A,B and C are just fixed numbers and they are whole numbers

OpenStudy (anonymous):

4x+2y=0 then?

jimthompson5910 (jim_thompson5910):

Example 7x + 10y = 13 is in standard form (this isn't the answer)

jimthompson5910 (jim_thompson5910):

y=-2x+4 is not in standard form, but you can do one move to get it into standard form

jimthompson5910 (jim_thompson5910):

notice the position of the x terms Ax+By = C y=-2x+4

jimthompson5910 (jim_thompson5910):

you need to get y=-2x+4 into the form Ax+By = C

OpenStudy (anonymous):

how do you do that

jimthompson5910 (jim_thompson5910):

notice how x is on the left side in Ax+By = C but x is on the right side in y=-2x+4

OpenStudy (anonymous):

so you have to flip it?

jimthompson5910 (jim_thompson5910):

how would you move that "2x" over

jimthompson5910 (jim_thompson5910):

what can be done to both sides

OpenStudy (anonymous):

i have no idea lol

jimthompson5910 (jim_thompson5910):

-2x+4 is the same as 4 - 2x, agreed?

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

so what operation undoes the subtraction of 2x

OpenStudy (anonymous):

the addition

jimthompson5910 (jim_thompson5910):

adding 2x to both sides, yes

jimthompson5910 (jim_thompson5910):